Sample problem: A particular sample held in a cell of 1 cm path length has an absorbency of 0.52 at 645nm, 1.02 at 663nm and 0.02 at 700m,. What are the concentrations of chlorophylls `a' and `b' in the sample? Do not forget to subtract the readings at 700nm from those at 645 and 663nms! (Answers are: 11.4ug/ml Chl a, 6.8ug/ml Chl b). I do not know how to get these answers.
1
Expert's answer
2020-09-21T06:31:14-0400

Molar mass of a-form: 893.51

A = ecl

l = 1

A = 1.02 - 0.02 = 1.00

e = 78,750

So c = 1/78,750 = 1.27x10-5 M

x 893.51 =

0.01135 g/L =

= 11.4 mg/mL =

= 11.4 ug/mL - Chl a

The concentration of b-form is calculated same way